OpenStudy (anonymous):

A cookie company packages its cookies in rectangular prism boxes designed with square bases that have both a length and a width of 4in. less than the height of the box. Which polynomial below represents the volume of a box with height x inches V=x(x+4)2 V=x(x−4)2 V=x(x−4) V=2x(x+4)

OpenStudy (anonymous):

Its B V=x(x−4)2 because x is the height, so x-4 is the length and also the width since length and a width of 4in. less than the height. So L = x-4, W = x-4, H = x Volume = length * width * height V = L * W * H V = (x-4)*(x-4)*x V=x(x−4)2

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

Welcome :)